OpenStudy (anonymous):

Triangle abc is ~ def ab=7x bc=4 de=7 ef=x what is the length of ab

OpenStudy (anonymous):

if the triangles are similar that means they are proportional right?

OpenStudy (anonymous):

when you set up the proportion problem 7x/4=7/x x=2, 7(2) ab=14